OpenStudy (anonymous):

-3/5 divided by -6/7 = A. 18/35 B. 7/10 C. 10/7 D. 35/18

OpenStudy (anonymous):

\[\frac{ \frac{ -3 }{ 5 } }{ \frac{ -6 }{ 7 } }\]First separate the problem so it can look more simpler.\[\frac{ -3 }{ 5 } \div \frac{ -6 }{ 7 }\]

OpenStudy (anonymous):

Then flip the second fraction so it can look like this:\[\frac{ -3 }{ 5 } \times \frac{ 7 }{ -6 }\] (*Remember when flipping the second fraction, the division sign changes to multiplication*)

OpenStudy (anonymous):

Now multiply the numerators together and the denominators. \[\frac{ -3 }{ 5 } \times \frac{ 7 }{ -6 } = \frac{ -21 }{ -30 }\]

OpenStudy (anonymous):

Lastly, you need to simplify the fraction.\[\frac{ -21 }{ -30 } = \frac{ 7 }{ 10 }\]

OpenStudy (anonymous):

Did you understand that?

OpenStudy (anonymous):

Thanks so much! Helped a lot!

OpenStudy (anonymous):

You're welcome!